An ambulance travels down a highway at 120kmh-1, its siren emitting sound of frequency 400Hz. What is the frequency heard by a passenger in a car traveling at 90kmh-1 as the car approaches and

by-passes the ambulance? 



1
Expert's answer
2021-03-08T08:32:57-0500

a) We can find the frequency from the Doppler Shift formula:


"f_o=f_s(\dfrac{v+v_o}{v-v_s}),""f_o=400\ Hz\cdot(\dfrac{340\ \dfrac{m}{s}+25\ \dfrac{m}{s}}{340\ \dfrac{m}{s}-33.33\ \dfrac{m}{s}})=476\ Hz."

(b) We can find the frequency from the Doppler Shift formula:


"f_o=f_s(\dfrac{v-v_o}{v+v_s}),""f_o=400\ Hz\cdot(\dfrac{340\ \dfrac{m}{s}-25\ \dfrac{m}{s}}{340\ \dfrac{m}{s}+33.33\ \dfrac{m}{s}})=337.5\ Hz."